please help me to understand substitution. here is an example 2x+5y=5
x+3y=4

Found 2 solutions by solver91311, lynnlo:
Answer by solver91311(24713)



Solve one of the equations for one of the variables in terms of the other. In your example, it is convenient to solve the second equation for in terms of because you only have to add to both sides.



Now that you have one of the variables defined in terms of an expression involving the other variable, replace (substitute) the defined variable in the other equation with the expression. In your example, replace in the first equation with



Solve the single variable equation that remains. Your example: solve for .

Once you have a value for one of the variables ( in your case) substitute that value back into either of the original equations and then solve for the remaining variable.

Report the solution as an ordered pair,

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lynnlo(4176)

2X+5Y=5
X+3Y=4
USING SUBSTITUTION METHOD:
X AS A FUNCTION OF Y(FROM THE FIRST EQUATION X=(5-5Y)/2
SUBSTITUTE INTO SECOND EQUATION
1(5+5Y)/2+3Y=4
SIMPLIFY:
1Y=3 OR Y=3
SUBSTITUTING Y=3 BACK INTO THE X EXPRESSED AS FUNCTION OF Y
X=(5-5Y)/2=-5